# Saddle Hierarchy in Dense Associative Memory

### v1

@@ 4.3 Fast training with splitting steepest descent
In Fig. (6), we compare DAM training time and accuracy with and without splitting steepest descent, using round markers and error bars to show their respective means and standard deviations as a function of the maximum number of hidden units $P_{\text{max}}$. Splitting stops at $P_{\text{max}}$ hidden units, and DAMs without splitting have $P_{\text{max}}$ hidden units from the start. All hidden units are split in each while-loop iteration of Alg. (1), except for the last iteration, where exactly $P_{\text{max}}-P_{\text{cur}}$ hidden units are split. As wanted, DAMs have similar accuracy with and without splitting (left panel). Moreover, the relatively small error bars of the training time indicate that the residual noise of the background processes is small and that the seed has a limited impact on the training speed (right panel). Splitting steepest descent has a significant speed advantage that scales very advantageously with $P_{\text{max}}$ (right panel). Without splitting, the training time is proportional to $P_{\text{max}}$. With splitting, the training time is almost constant. The clearest sign that it increases with $P_{\text{max}}$ is the small jump between $P_{\text{max}}=1500$ and $P_{\text{max}}=1625$, where the number of while-loop iterations in Alg. (1) increases by 1. In other words, the run time of Alg. (1) is better explained as proportional to $\left\lceil\log P_{\text{max}}\right\rceil$ than $P_{\text{max}}$ in the range of $P_{\text{max}}$ of our numerical experiment, which is a dramatic improvement over the run time without splitting.
Figure 6: The classification accuracy and training time of dense associative memory (DAM) networks trained with and without splitting as a function of the maximum number of hidden unit $P_{\text{max}}$. The round markers and their corresponding error bars show the means and standard deviations of the measurements made at each $P_{\text{max}}$. Statistics are collected from 1010 different random seeds. In the left panel, we verify that the DAMs with and without splitting have a similar accuracy. In the right panel, we compare their training times $t$. To facilitate this comparison, we normalize the $t$-axis such that the data point indicated by the arrow is at $t=1$.
@@ 5 Conclusion
In this work, we study a dense associative memory (DAM) model capable of learning interpretable solutions to both supervised and unsupervised problems, which we construct from the framework of Boltzmann machines.
@@
We derive two sets of equations that respectively characterize the stationary points of DAMs trained on real data using maximum likelihood estimation and the fixed points of DAMs trained on synthetic data in the teacher-student setting. Guided by their similarity, we then establish that the maximum likelihood equations are a special case of the teacher-student equations. Building on this equivalence, we introduce an effective loss function that significantly stabilizes training thanks to a regularization parameter that mimics the noise present in the data generation process.
@@
We show that the stationary points of the effective loss of DAMs with a given number of hidden units are saddle points of larger DAMs. We considerably accelerate training using a custom implementation of the splitting steepest descent network-growing algorithm [36, 37] inspired by this saddle-point hierarchy principle. In our numerical experiments, the training time of splitting steepest descent is approximately logarithmic in the number of hidden units, which is a huge improvement over the linear training time that we observe without splitting steepest descent. Further research is needed to clearly determine the asymptotic functional form of the training time.
@@
In Section 4.1, we briefly discuss how our DAM is interpretable. The ingredients needed to make an interpretable neural network are not well understood, and we believe that our DAM offers a promising opportunity to explore that area. Studying the teacher-student setting with teacher inverse temperature $\beta^{*}$ less than order $N$ is another interesting research direction that we leave to future work. Beyond the framework of our model, [98, 99] have designed a technique that trains Gaussian mixture models by replicating mixture components. In contrast with our work, it uses a handpicked cooling schedule and breaks permutation symmetries with random noise. By design, splitting steepest descent finds the directions in the loss landscape that allow NNs to escape saddle points with permutation symmetries as quickly as possible, but finding these directions is more expensive than generating random noise (see Appendix H and [36, 37]). It would be interesting to study this tradeoff and identify the regimes where each method is preferable over the other. More generally, one could investigate how permutation symmetry-breaking takes place in energy models related to DAMs, such as attentional BMs [27], and study its relationship with the permutation symmetry-breaking [81, 84] and dynamical transitions [70, 71, 100, 73, 101] found in other types of RBMs. [11, 102] have made connections between DAM learning dynamics and cellular differentiation. Similarly, it could be possible to investigate whether there are similarities between cellular division and splitting steepest descent in our model. Finally, it would be interesting to see if splitting steepest descent and our proposed regularization technique can be used to improve the training of transformers and generative diffusion.
@@
The code for the numerical experiments presented in this work is available at this public repository [38].
@@ Appendix A Derivation of the model
where $\mathbf{J}=\left\{\mathbf{w},\mathbf{u},\mathbf{b}\right\}$ is the set of all RBM parameters, i.e. $\mathbf{w}=\left\{w^{\mu}_{i}\right\}_{1\leq i\leq N}^{1\leq\mu\leq P}$, $\mathbf{u}=\left\{u^{\mu}_{i}\right\}_{1\leq i\leq C}^{1\leq\mu\leq P}$ and $\mathbf{b}=\left\{b^{\mu}\right\}_{\mu=1}^{P}$. The prior densities $\operatorname{P}_{0}\left(\mathbf{h}\right)$ and $\operatorname{P}_{0}\left(\mathbf{q}\right)$ are non-zero only when $\mathbf{h}\in\left\{\mathbf{e}_{\gamma}\right\}_{\gamma=0}^{P}$ and $\mathbf{q}\in\left\{\mathbf{e}_{y}\right\}_{y=0}^{C}$, in which case we have
@@
where $Z_{\beta}\left(\mathbf{J}\right)$ is an unknown normalization constant. The data $\mathbf{x}$ belonging to each cluster $\mu>0$ follows a von Mises-Fisher (vMF) distribution $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)\propto\exp\left(\beta\sum_{i=1}^{N}w^{\mu}_{i}x_{i}\right)$ centered on $\mathbf{w}^{\mu}=\left\{w^{\mu}_{i}\right\}_{i=1}^{N}$ (see Appendix C). We assume that $\sqrt{\sum_{i=1}^{N}\left(w^{\mu}_{i}\right)^{2}}=1$ so that the corresponding cluster centroids $\mathbf{w}^{\mu}$ lie on the hypersphere $S^{N-1}$ like the data $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{N}$. We marginalize $\operatorname{P}_{\beta}\left(\mathbf{x},y,\mu|\mathbf{J}\right)$ over the hidden units $\mu$ and get
@@
Appendix B Stationarity conditions of the loss
$\sum_{i=1}^{N}\left[w^{\mu}_{i}\right]^{2}=1$
In Appendix F, we show that its solution is
for all $0\leq\gamma\leq P$, where the normalization constant $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ is defined using Eqs. (F) of Appendix F. Combining these equations with the ones for $\mathbf{w}$, we find the stationarity conditions
for all $1\leq\mu\leq P$ and $0\leq\gamma\leq P$.
Appendix C Integration of the von Mises-Fisher density
@@
Appendix D Replicated partition function and free entropy

### v2

@@ 4.3 Fast training with splitting steepest descent
In Fig. (7), we compare DAM training time and accuracy with and without splitting steepest descent, using round markers and error bars to show their respective means and standard deviations as a function of the maximum number of hidden units $P_{\text{max}}$. Splitting stops at $P_{\text{max}}$ hidden units, and DAMs without splitting have $P_{\text{max}}$ hidden units from the start. All hidden units are split in each while-loop iteration of Alg. (1), except for the last iteration, where exactly $P_{\text{max}}-P_{\text{cur}}$ hidden units are split. As wanted, DAMs have similar accuracy with and without splitting (left panels). Moreover, the reasonably small error bars of the training time indicate that the residual noise of the background processes is controlled and that the seed has a limited impact on the training speed (right panels). Splitting steepest descent has a significant speed advantage that scales very advantageously with $P_{\text{max}}$ (right panels). Without splitting, the training time is proportional to $P_{\text{max}}$. With splitting, it appears to be piecewise constant. The clearest sign that it increases with $P_{\text{max}}$ is the jump between $P_{\text{max}}=1500$ and $P_{\text{max}}=1625$, where the number of while-loop iterations in Alg. (1) increases by 1. In other words, the run time of Alg. (1) is better explained as proportional to $\left\lceil\log P_{\text{max}}\right\rceil$ than $P_{\text{max}}$ in the range of $P_{\text{max}}$ of our numerical experiment, which is a dramatic improvement over the run time without splitting. The benefit of the algorithm becomes evident primarily when the accuracy improves consistently across a wide range of hidden unit counts, i.e. when training a substantially large network is beneficial. This pattern, for example, does not apply to Fashion MNIST (see Fig. 7, bottom plot), where the accuracy gain from training a larger network is marginal. As a result, the need for strategies that accelerate training becomes less critical, although such acceleration still occurs.
Figure 7: The classification accuracy and training time of dense associative memory (DAM) networks trained on MNIST [50] and Fashion-MNIST [77] with and without splitting as a function of the maximum number of hidden unit $P_{\text{max}}$. The round markers and their corresponding error bars show the means and standard deviations of the measurements made at each $P_{\text{max}}$. Statistics are collected from 1010 different random seeds. In the left panels, we verify that the DAMs with and without splitting have a similar accuracy. In the right panels, we compare their training times $t$. To facilitate this comparison, we normalize the $t$-axis such that the data point indicated by the arrow is at $t=1$.
@@ 5 Conclusion
In this work, we study a Dense Associative Memory model based on the framework of Boltzmann machines that is capable of learning interpretable solutions to both supervised and unsupervised problems.
@@
We derive two sets of equations that respectively characterize the stationary points of DAMs trained on real data using maximum likelihood estimation and the fixed points of DAMs trained on synthetic data in the teacher-student setting. Guided by their similarity, we then establish that the maximum likelihood equations are a special case of the teacher-student equations. Building on this equivalence, we introduce an effective loss function that makes training significantly more stable thanks to a regularization parameter that mimics the noise present in the data generation process.
@@
We show that the stationary points of the effective loss of DAMs with a given number of hidden units are saddle points of larger DAMs, and we considerably accelerate training using a custom implementation of the splitting steepest descent network-growing algorithm [29] inspired by this saddle-point hierarchy principle. In our numerical experiments, the training time of splitting steepest descent is approximately logarithmic in the number of hidden units, which is a huge improvement over the linear training time that we observe without splitting steepest descent. Further research is needed to clearly determine the asymptotic functional form of the training time.
@@
In Section 4.1, we briefly discuss how our DAM is interpretable. The ingredients needed to make an interpretable neural network are not well understood, and we believe that our DAM offers a promising opportunity to explore that area. Studying the teacher-student setting with teacher inverse temperature $\beta^{*}$ less than order $N$ is another interesting research direction that we leave to future work. Beyond the framework of our model, [78, 79] have designed a technique that trains Gaussian mixture models by replicating mixture components. In contrast with our work, it uses a handpicked cooling schedule and breaks permutation symmetries with random noise. By design, splitting steepest descent finds the directions in the loss landscape that allow NNs to escape saddle points with permutation symmetries as quickly as possible, but finding these directions (as described in Appendix K and [29, 75]) is more expensive than generating random noise. It would be interesting to study this tradeoff and identify the regimes where each method is preferable over the other. More generally, one could investigate how permutation symmetry breaking takes place in energy models related to DAMs, such as attentional BMs [80], and study its relationship with the permutation symmetry-breaking [62, 63] and dynamical transitions [53, 54, 81] found in other types of RBMs, whose careful analysis was also used to improve training [82]. Recent works [24, 83] have made connections between DAM learning dynamics and cellular differentiation. Similarly, it could be possible to investigate whether there are similarities between cellular division and splitting steepest descent in our model. Finally, it would be interesting to see if splitting steepest descent and our proposed regularization technique can be used to improve the training of transformers and generative diffusion.
@@
The code for the numerical experiments presented in this work is available in the following public repository [30].
@@ Appendix A Derivation of the model
where $\mathbf{J}=\left\{\mathbf{w},\mathbf{u},\mathbf{b}\right\}$ is the set of all RBM parameters, i.e. $\mathbf{w}=\left\{w^{\mu}_{i}\right\}_{1\leq i\leq N}^{1\leq\mu\leq P}$, $\mathbf{u}=\left\{u^{\mu}_{i}\right\}_{1\leq i\leq C}^{1\leq\mu\leq P}$ and $\mathbf{b}=\left\{b^{\mu}\right\}_{\mu=1}^{P}$. The prior densities $\operatorname{P}_{0}\left(\mathbf{h}\right)$ and $\operatorname{P}_{0}\left(\mathbf{q}\right)$ are nonzero only when $\mathbf{h}\in\left\{\mathbf{e}_{\gamma}\right\}_{\gamma=0}^{P}$ and $\mathbf{q}\in\left\{\mathbf{e}_{y}\right\}_{y=0}^{C}$, in which case we have
@@
where $Z_{\beta}\left(\mathbf{J}\right)$ is an unknown normalization constant. The data $\mathbf{x}$ belonging to each cluster $\mu>0$ follows a von Mises-Fisher (vMF) distribution $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)\propto\exp\left(\beta\sum_{i=1}^{N}w^{\mu}_{i}x_{i}\right)$ centered on $\mathbf{w}^{\mu}=\left\{w^{\mu}_{i}\right\}_{i=1}^{N}$ (see Appendix B). We assume that $\sqrt{\sum_{i=1}^{N}\left(w^{\mu}_{i}\right)^{2}}=1$ so that the corresponding cluster centroids $\mathbf{w}^{\mu}$ lie on the hypersphere $S^{N-1}$ like the data $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{N}$. We marginalize $\operatorname{P}_{\beta}\left(\mathbf{x},y,\mu|\mathbf{J}\right)$ over the hidden units $\mu$ and get
@@
Appendix B Integration of the von Mises-Fisher density
@@
Appendix C Normalization of the weights
$\sum_{\gamma=0}^{P}p^{\gamma}_{y}$ $=p_{\mathbf{q}}\left(y\right)$ (20)
$\sum_{y=0}^{C}p^{\gamma}_{y}$ $=p_{\mathbf{h}}\left(\gamma\right).$
$\frac{\bar{p}^{\gamma}_{y}}{p^{\gamma}_{y}}$ $=\lambda_{y}+\omega^{\gamma},$
$p^{\gamma}_{y}$ $=\frac{1}{\lambda_{y}+\omega^{\gamma}}\bar{p}^{\gamma}_{y}.$
Using the row and column constraints $\sum_{\gamma=0}^{P}p^{\gamma}_{y}=p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}p^{\gamma}_{y}=p_{\mathbf{h}}\left(\gamma\right)$, we find that the Lagrange multipliers $\lambda_{y}$ and $\omega^{\gamma}$ solve the nonlinear equations
For conciseness, we define $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)=\lambda_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)+\omega^{\gamma}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$, where $\lambda_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ and $\omega^{\gamma}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ are the $\lambda_{y}$ and $\omega^{\gamma}$ solving Eqs. (C) at given $P$ and $p_{\mathbf{h}}$. Using these definitions, we find the implicit solution $p^{\gamma}_{y}=\bar{p}^{\gamma}_{y}/\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$. As wanted, this equation is useful in analytical calculations involving the saddle-point equations and the stationarity conditions of the loss. However, it is also quite slow to solve numerically for a given $\mathbf{\bar{p}}$, as reported in multiple studies [90, 91, 92]. Therefore, although we can train $p^{\gamma}_{y}$ by iterating $p^{\gamma}_{y}=\bar{p}^{\gamma}_{y}/\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$, it is not a very efficient method.
In order to efficiently train $p^{\gamma}_{y}$, we devise a faster way to solve Eqs. (C) than through Eq. (C). Exponentiating both sides of the first line of Eqs. (C), we find
where $\eta$ is an arbitrary scalar that will play the role of a learning rate. Fix $k^{\gamma}_{y}=p^{\gamma}_{y}\exp\left(\eta\partial_{p^{\gamma}_{y}}f\left(\mathbf{p}\right)\right)$. By the Sinkhorn-Knopp theorem [93, 94], there is a rescaled matrix of the form $p^{\prime\gamma}_{y}=D^{\gamma}_{L}\left(\mathbf{k}\right)k^{\gamma}_{y}D^{y}_{R}\left(\mathbf{k}\right)$ that satisfies the same constraints $p^{\prime\gamma}_{y}\geq 0$, $\sum_{\gamma=0}^{P}p^{\prime\gamma}_{y}=p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}p^{\prime\gamma}_{y}=p_{\mathbf{h}}\left(\gamma\right)$ as $p^{\gamma}_{y}$ if some technical conditions are satisfied [95]. Moreover, if $p^{\prime\gamma}_{y}$ exists, then it is unique [96, 97], and we can quickly compute suitable scaling factors $D^{\gamma}_{L}\left(\mathbf{k}\right)$ and $D^{\gamma}_{R}\left(\mathbf{k}\right)$ for $\mathbf{k}=\left\{k^{\gamma}_{y}\right\}^{0\leq\gamma\leq P}_{0\leq y\leq C}$ using the Sinkhorn-Knopp algorithm first proposed in [98] and theoretically justified in subsequent works [93, 94]. See [95] for a concise review of all these points. $p^{\prime\gamma}_{y}$ is generally not equal to $p^{\gamma}_{y}$. However, we observe that the iteration
converges at small $\eta$ (for example $\sim 0.1/P$), which means that it must converge to a solution of Eqs. (C). Since the Sinkhorn-Knopp algorithm is fast, it is significantly more efficient to iterate Eqs. (C) than $p^{\gamma}_{y}=\bar{p}^{\gamma}_{y}/\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ to solve Eqs. (C).
In practice, we train $p^{\gamma}_{y}$ using a stochastic variant of Eqs. (C) where the gradient is estimated over small batches of data and smoothed with a momentum hyperparameter. Furthermore, we multiply the gradient by $\delta_{yy^{\prime}}-p^{\gamma}_{y}/p_{\mathbf{h}}\left(\gamma\right)$ to reduce the size of the components of $\exp\left(\eta\partial_{p^{\gamma}_{y}\left(t\right)}f\left(\mathbf{p}\left(t\right)\right)\right)$ that move $p^{\gamma}_{y}$ away from the constraints $\sum_{\gamma=0}^{P}p^{\gamma}_{y}=p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}p^{\gamma}_{y}=p_{\mathbf{h}}\left(\gamma\right)$. This step is a simple approximation of the gradient projection proposed in [99].
Appendix D Initialization and learning rate
To make all our figures, except Fig. (6), we initialize the memories $\mathbf{w}^{\mu}$ uniformly at random on the unit hypersphere $S^{N-1}$ [100]. To make Fig. 6, we instead use the algorithm of [101] to sample the initial memories $\mathbf{w}^{\mu}$ from a vMF distribution (see Appendix B) with mean direction $\tilde{\mathbf{x}}^{*}=\bar{\mathbf{x}}^{*}/\|\bar{\mathbf{x}}^{*}\|$, where $\bar{\mathbf{x}}^{*}=\frac{1}{N}\sum_{\mu_{*}=1}^{P^{*}}\mathbf{x}^{*\mu_{*}}$ is the mean of the patterns $\mathbf{x}^{*\mu_{*}}$ and $\|\bar{\mathbf{x}}^{*}\|=\sqrt{\sum_{i=1}^{N}\left[\bar{x}^{*}_{i}\right]^{2}}$. By construction, the mean direction $\tilde{\mathbf{x}}^{*}$ is the ordered solution of Eq. (3.1) with the most permutation symmetries, or in other words the root of the learning dynamics tree shown in Fig. 6, so initializing the memories around it helps reveal the tree structure.
We use a learning rate of $\eta\sim 0.1$ to train the memories $\mathbf{w}^{\mu}$. Based on our experience, $\mathbf{p}$ trains well when its own learning rate is approximately $\left(1+\frac{p_{\mathbf{h}}\left(0\right)}{1-p_{\mathbf{h}}\left(0\right)}\right)P$ times smaller. Without this rescaling, the multiplicative update factor $\exp\left(\eta\partial_{p^{\gamma}_{y}}L\left(\mathbf{w},\mathbf{p}\right)\right)$ described in Appendix C can be relatively large even when $\eta$ is relatively small, making it ill behaved. Equivalently, we can also train $g^{\gamma}_{y}=\left(1+\frac{p_{\mathbf{h}}\left(0\right)}{1-p_{\mathbf{h}}\left(0\right)}\right)P\mathbf{p}$ with the same learning rate $\eta$ as $\mathbf{w}$. We adopt this approach in our code available at [30].
[…]
Algorithm 2 Rescaled splitting steepest descent [29, 75]
3:min $L\left(\mathbf{w},\mathbf{g}\right)$ with SGD
4:while $P_{\text{cur}}<P_{\text{max}}$ do
11:  Update $P_{\text{cur}}\leftarrow P_{\text{cur}}+R$
13:  min $L\left(\mathbf{w},\mathbf{g}\right)$ with SGD
14:end while
15:return $\triangleright$ See Appendix K and [29, 75] for details about steps 5 and 12
When we train our DAM without splitting steepest descent, we set $P_{0}=1$. As discussed in [102, 103], the resulting weights are, in some sense, the most “typical” for the constraints $g^{\gamma}_{y}\geq 0$, $\sum_{\gamma=0}^{P}g^{\gamma}_{y}=\left(P+1\right)p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}g^{\gamma}_{y}=\left(P+1\right)p_{\mathbf{h}}\left(\gamma\right)=1$ (see Section 2). With splitting steepest descent, we set $P_{0}=P_{\text{cur}}/P_{\text{final}}$ to ensure that $g_{\mathbf{h}}\left(0\right)=\sum_{y=0}^{C}g^{0}_{y}$ is approximately the same size as the other entries of $g_{\mathbf{h}}\left(\gamma\right)$ after training. For simplicity, we set $p_{\mathbf{q}}\left(y\right)$ to the proportions of classes in the data during supervised training. For unsupervised training (see Section 4.2), we break the permutation symmetry between the different classes by using different values for all the entries of $p_{\mathbf{q}}\left(y\right)$. Otherwise, class weights with unbroken permutation symmetries stay stuck at their initial conditions.
Appendix E Important symbols introduced in the Model Section
The following table summarizes the important symbols introduced in Section 2 and their meaning.
Symbol Meaning
$\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{N}$ Data layer
$\mathbf{h}=\left\{h_{\mu}\right\}_{\mu=1}^{P}$ Hidden layer
$\mathbf{q}=\left\{q_{y}\right\}_{y=1}^{C}$ Class layer
NN # data units
$P$ # hidden units (and clusters)
$C$ # class units (and classes)
$\mathbf{J}$ Set of all trainable weights
$\mathbf{w}=\left\{w_{i}^{\mu}\right\}_{1\leq i\leq N}^{1\leq\mu\leq P}$ Weight matrix of the data layer
$\mathbf{p}=\left\{p^{\gamma}_{y}\right\}_{0\leq y\leq C}^{0\leq\gamma\leq P}$ Weight matrix of the class layer
$p_{\mathbf{q}}\left(y\right)$ Marginal distribution $\sum_{\gamma=0}^{P}p^{\gamma}_{y}$
$p_{\mathbf{h}}\left(\gamma\right)$ Marginal distribution $\sum_{y=0}^{C}p^{\gamma}_{y}$
$\beta$ Inverse temperature
$\Omega_{N}\left(\beta\right)$ Normalization constant of the vMF distribution
$\mathbf{x}^{*\mu}=\left\{x^{*\mu}_{i}\right\}_{i=1}^{N}$ Pattern from a given dataset
$\mathbf{q}^{*\mu}=\left\{q^{*\mu}_{y}\right\}_{y=0}^{C}$ Soft label of the pattern
$P^{*}$ # hidden units (and clusters) of the teacher
$\mathbf{w}^{*}=\left\{w_{i}^{*\mu}\right\}_{1\leq i\leq N}^{1\leq\mu\leq P}$ Weight matrix of the data layer of the teacher DAM
$\mathbf{p}^{*}=\left\{p^{*\gamma}_{y}\right\}_{0\leq y\leq C}^{0\leq\gamma\leq P}$ Weight matrix of the class layer of the teacher DAM
$\mathbf{g}^{*}$ Rescaled weight matrix $\mathbf{g}^{*}=P^{*}\mathbf{p}^{*}$ of the class layer of the teacher DAM
$p_{\mathbf{q}}^{*}\left(y\right)$ Marginal distribution $\sum_{\gamma=0}^{P}p^{*\gamma}_{y}$
$p_{\mathbf{h}}^{*}\left(\gamma\right)$ Marginal distribution $\sum_{y=0}^{C}p^{*\gamma}_{y}$
$\beta^{*}$ Inverse temperature of the teacher
$\mathcal{D}=\left\{\mathbf{x}^{c},y^{c}\right\}_{c=1}^{M}$ Set of data points $\mathbf{x}^{c}$ and labels $y^{c}$ generated by the teacher
$M$ Number of data points generated by the teacher
$\alpha=M/N$ Load of data generated by the teacher
Table 2: Summary of the important symbols introduced in Section 2 and their meaning.
Appendix F Stationarity conditions of the loss
$\sum_{i=1}^{N}\left[w^{\mu}_{i}\right]^{2}=1$
In Appendix C, we show that its solution is
for all $0\leq\gamma\leq P$, where the normalization constant $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ is defined using Eqs. (C) of Appendix C. Combining these equations with the ones for $\mathbf{w}$, we find the stationarity conditions
for all $1\leq\mu\leq P$ and $0\leq\gamma\leq P$.
Appendix G Replicated partition function and free entropy